OpenStudy (anonymous):

solve the equation for x: 1 - sin(x) = (root3)(cos(x)) 2sin^3(x)+sin^2(x)-2sinx=1

OpenStudy (anonymous):

just a start would be incredible

OpenStudy (anonymous):

are these two separate problems or a system of equations?

OpenStudy (anonymous):

two different problems, forgot the s right after solve the equation, sorry!

OpenStudy (anonymous):

I'll help you with the first problem and you can post the second as a different question. By root do you mean square root? (sqrt)

OpenStudy (anonymous):

I think i understand the first problem

OpenStudy (anonymous):

its the double angle formula that is required, and after that it's pretty straightforward

OpenStudy (slaaibak):

square both sides in the first problem: (1 - sin(x))^2 = ((root3)(cos(x)))^2 1 - 2sinx + sin^2 x = 3 cos^2 x 1 - 2sinx + sin^2x = 3(1-sin^2 x) now it should be easy from here

OpenStudy (anonymous):

for my solution set I got: \[x = \pi/6 + 2\pi n\]\[x=3\pi/2 + 2\pi n\]

OpenStudy (slaaibak):

4x^2 - 2sinx -2 = 0 2sin^2 - sinx - 1 = 0 (2sinx + 1)(sin x - 1) = 0 sin x = -1/2 sin x = 1 so solution set: x = -pi/6 + 2npi or x = pi/2 + 2npi

OpenStudy (anonymous):

doesn't 7pi/6 also make sinx=-1/2?

OpenStudy (anonymous):

So then there would be three solution sets if what I said above is true.

OpenStudy (slaaibak):

because we squared the equations, all the solution sets won't work, because some will give a positive RHS while the LHS gives a negative, and vice versa. Consider 7pi/6: LHS: 1 - (-1/2) = 3/2 RHS: sqrt3 * -sqrt3/2 = -3/2 Therefore RHS and LHS are not equal

OpenStudy (anonymous):

I see, so then it is only be pi/2 and 11pi/6?

OpenStudy (slaaibak):

if the solution needs to be in [0, 2pi], yes

OpenStudy (anonymous):

awesome! So for the second problem, I took out the sinx and made it 2sin^x+sinx-2, all of that being equal to 1. Does that sound like the right way to start it?

OpenStudy (slaaibak):

2sin^3(x)+sin^2(x)-2sinx=1 for this one, it's important to realize you can use a group factor method: Rewrite it like this: 2sin^3(x)+sin^2(x)-2sinx -1 = 0 you can see absolute value of the coefficients, in order of powers (rank), goes: 2 1 2 1 Now you can do this: sin^2 x(2sin x + 1) - (2sin x + 1) = 0 -> (2sin x + 1)(sin^2 x - 1) = (2sin x + 1)(sin x + 1)(sin x - 1) = 0 now you can solve from there

OpenStudy (anonymous):

I found 4 solution sets: \[3\pi/2 + 2\pi n\]\[\pi/2 + 2\pi n\]\[11\pi/6+2\pi n\]\[7\pi/6 + 2\pi n\]

OpenStudy (anonymous):

Checked them all as well, none are extraneous